An object is thrown vertically upwards with initial speed 15m/s. After 1s, a second object is thrown up. What is the initial speed that the second object must have to hit the first object at a height of 11m?
Answer #1

12 Nd shab 46 conno =) 42(t-1) = 11+4.9 (t-1)2 96 t = 1.2177 sec. 42 = 11+ 4.9+(1.2177-1) 2 ა — (1.2177-1) ab Now coord 142=5